Suppose the mean number of days to germination of a variety of seeds is 34, with a standard deviation of 4.3 days. What is the probability that the mean germination time of a sample of 250 seeds will be within 0.5 days of the population mean?




1
Expert's answer
2022-03-15T03:20:15-0400

"\sigma=4.3,\mu=34,n=250"

Standard deviation of sampling distribution "=\frac{sd}{\sqrt{n}}=\frac{4.3}{\sqrt{250}}=0.2720"

"P(\frac{-0.5}{0.2720}<Z<\frac{0.5}{0.2720})=P(-1.84<Z<1.84)"

"=P(Z<1.84)-P(Z<-1.84)=0.9671-0.0329"

"=0.9342"